A Science Fiction Story Problem: Many Years in the future (2166 A.D.) the Mexican army builds a great wall (see next page forSeveral students discussing the circuit below where bulbs C and B are brighter than A and D, make the following contentions.
Answer #1

We will agree with Arthur. As there is only one path, so the rate of flow of charge i.e. current flowing through the circuit should be same every where. Consider any point in the circuit, if any electron comes towards that, same electron should leave as well, there cannot be any accumulation of electrons at any point, it is in accordance with the law of conservation of charge. As the bulbs C and B are brighter than A and D, their ohmic resistance must be more than that of A and D. If I is the current in the circuit, the power dissipated in each bulb is given by the formula, P = I^2*R. So higher the resistance, more the power dissipated in that bulb and higher will be its brightness as the current is same.

Let us suppose 0.1 A is the current flowing in the circuit. Resistance of B and C is 2\Omega each and that of A and D 1\Omega each. Then power dissipated in both B and C is P1=(0.1)^2*2 watt = 0.02 W each and in both A and D it is P2 = (0.1)^2*1 W=0.01 W each. Therefore we can see that P1 > P2 and hence B and C will glow brighter.

Potential differences across both A and D are same as potential difference is given by the relation, V = I*R, where R is the resistance and I is the current. As both current and resistance are same for both A and D, hence there is same potential difference across each one of them.